# Learning UAV-based path planning for efficient localization of objects using prior knowledge

### v1

@@ Abstract
UAV’s are becoming popular for various object search applications in agriculture, however they usually use time-consuming row-by-row flight paths. This paper presents a deep-reinforcement-learning method for path planning to efficiently localize objects of interest using UAVs with a minimal flight-path length. The method uses some global prior knowledge with uncertain object locations and limited resolution in combination with a local object map created using the output of an object detection network. The search policy could be learned using deep Q-learning. We trained the agent in simulation, allowing thorough evaluation of the object distribution, typical errors in the perception system and prior knowledge, and different stopping criteria. When objects were non-uniformly distributed over the field, the agent found the objects quicker than a row-by-row flight path, showing that it learns to exploit the distribution of objects. Detection errors and quality of prior knowledge had only minor effect on the performance, indicating that the learned search policy was robust to errors in the perception system and did not need detailed prior knowledge. Without prior knowledge, the learned policy was still comparable in performance to a row-by-row flight path. Finally, we demonstrated that it is possible to learn the appropriate moment to end the search task. The applicability of the approach for object search on a real drone was comprehensively discussed and evaluated. Overall, we conclude that the learned search policy increased the efficiency of finding objects using a UAV, and can be applied in real-world conditions when the specified assumptions are met.
@@ 1 Introduction
In recent years, Unmanned Aerial Vehicles (UAVs) are becoming increasingly popular for various applications in agriculture (Rejeb et al., 2022). Examples in agriculture are finding weeds or diseased plants in large arable fields (Albani et al., 2019; Chin et al., 2023), detecting cattle in pastures (Rivas et al., 2018; Liu et al., 2021) and blossom detection in orchards (Zhang et al., 2023). Generally, these applications have a main task: to find objects of interest in a area larger than the field-of-view (FoV) of the UAV, which requires the UAV to fly over the field. A limitation of using UAVs for these applications is their limited battery capacity (Rejeb et al., 2022; Gugan and Haque, 2023). Therefore it is important to find a path between all the objects of interest that minimizes the UAV’s flight time in order to increase the area that can be inspected in a single flight.
@@
When these objects are uniformly distributed over the area, a coverage path planner covering the entire search area is suitable and efficient. However, in applications where the objects are non-uniformly distributed, it may be more efficient to use a search policy that searches for objects instead of covering the whole area. A task such as the detection of weeds is an example of an application with non-uniformly distributed objects, since some weed species occur in distinct patches in a field (Cardina et al., 1997; Xu et al., 2023a). A lot of work is done on weed detection using drone images (Xu et al., 2023b; Anul Haq, 2022; Pei et al., 2022), however they all follow predefined, row-by-row, flight paths.
@@
Over the last years, Reinforcement Learning (RL) has gained more attention in path planning for both mobile robots (Yu et al., 2020; Gao et al., 2020; Niroui et al., 2019) as well as UAVs (Azar et al., 2021; Tu and Juang, 2023). These RL-based methods can learn search strategies through interaction with the environment by maximizing the information gathered in an environment (Lodel et al., 2022). RL offers several advantages for path planning: it does not need a detailed map of obstacles in the environment, can deal with noisy sensor information (Gao et al., 2020) and can learn spatial relations between objects of interest in the environment. When trained, a RL learned policy makes sequential decisions about the direction of the robot or drone with relatively short calculation times, as compared to traditional methods like Dijkstra’s algorithm, A* and D* (Tu and Juang, 2023). This allows for online path planning, eliminating the need to calculate the complete path in advance. Using RL it is possible to plan paths with less or uncertain prior knowledge about the environment (Yu et al., 2020; Gugan and Haque, 2023), and thereby making path planning more flexible and reactive to a changing environment.
Several works have studied path planning using RL. Panov et al. (2018) and Yu et al. (2020) showed that a trained RL network is able to generate a policy to move an agent in a grid-based world to a target location while avoiding obstacles. A novel neural network architecture for coverage path-planning and object search was introduced by Theile et al. (2020, 2021), combining detailed high-resolution local information with low resolution global information. However, these approaches assume full prior knowledge about the location of the target objects and the obstacles, which in many applications is uncertain or even absent. As example, the location of weeds in a field may not be exactly known beforehand, however, some prior knowledge can be derived from the locations of the weeds during the previous year.
When having full prior knowledge, there is enough information to stop the search task when all objects are found. However, when the prior knowledge becomes more sparse and unreliable, the number of objects might be unknown. This makes the end of the search task undefined. Yang et al. (2018) and Druon et al. (2020) used a learned stopping action, where the agent could decide to stop searching when it was not profitable anymore.
Training a RL agent requires huge amounts of data as it needs to explore an enormous state-action space (Gugan and Haque, 2023). To train a RL agent to localize objects in a field, we need the agent to collect a very large amount of training data containing a large variety in object locations in the field and allowing it to explore different policies. Because this is infeasible to achieve in the real world, we trained the agent using simulations, as is commonly done in literature, e.g. Azar et al. (2021); Gao et al. (2020); Theile et al. (2020, 2021); Panov et al. (2018); Yu et al. (2020). We ensured the simulation was realistic, as explained in the next paragraph. An additional benefit of a simulation is that we can investigate the effect of different object distributions, detection inaccuracies and prior knowledge uncertainty in greater detail and with a much higher number of repetitions than would be feasible in real-world experiments and thereby shows the practical potential of such RL based path planner for localizing objects of interest in a field.
@@
In this paper, we propose an RL-based path planner that works on a higher abstraction level, getting a local object map extracted from a perception module based on a camera image, and providing actions in terms of flight directions, that would be executed by a drone’s flight controller. Figure 1 illustrates how the resulting simulation-trained policy can be used in the real-world to control the drone. The drone’s camera image, containing variations in the object appearance due to, e.g., lightning conditions and natural variation, is converted to a local object map using an object detection network. The resulting object coordinates are placed in an object map, which is then used as input to the RL policy. The policy determines the action with respect to the highest expected reward. The discrete output actions (fly north, east, west, south) are then translated to motor commands, by the drone’s flight controller. This abstraction allows us to train the RL agent in simulation.
@@
Figure 1: High-level drone control using Reinforcement Learning (RL) in combination with a local object map from a perception system and some potential object locations from the prior knowledge. The output of the detection network is used together with prior knowledge as input for the RL policy, the resulting discrete actions are then translated by the drone’s flight controller to motor commands.
@@
It is important, however, that the abstract simulation includes typical errors in a perception model and a flight controller. The typical errors in translating the image into object coordinates are false positive and false negative detections and inaccuracies in the location of the detections. These are included in the simulation as explained in the materials section. Furthermore, the simulation makes some assumptions about the system: (a) it is assumed the drone is able to accurately execute discrete flight actions, such as fly one meter in a specific direction, (b) a maximum field size based on the number of grid-cells in the simulation and the area of each cell, (c) prior knowledge of the field shape and boundaries, (d) a trained detection network and (e) some potential locations of objects of interest.
@@
The objective of this paper is to develop and evaluate a RL-based method for path planning to efficiently localize objects of interest using UAVs with a minimal path length. Despite being a simulation study, we provide an important contribution to a real-world application by analyzing the effects of the spatial distribution of the objects of interest, typical errors in detection network output and of the typical errors associated with prior knowledge. The real world application is discussed at length in section 4.5. Specifically, we study (1) the effect of the distribution of the objects, (2) the effect of detection errors, (3) the effect of the quality of prior knowledge and finally, and (4) the effect of different stopping criteria.
@@ 2 Materials and Methods
In this chapter, we describe the simulation environment, problem statement, the used RL method and the experiments.
@@ 2.1 Simulation environment
The field is simulated as a squared grid of $M\times M\in\mathbb{N}^{2}$ grid cells, where $\mathbb{N}$ is the set of natural numbers. The real-world size of each grid cell depends on the required spatial resolution, which is not part of the simulation environment. In this field, the number of objects, $n$ is drawn from a normal distribution $\mathcal{N}_{\textrm{obj}}(\mu,\sigma)$. These objects are then distributed according to $k$ different multivariate Gaussian distributions, where $k$ is drawn from $\mathcal{N}_{\textrm{dist}}(\mu)$. Each Gaussian distribution $\mathcal{N}(\mu_{i},\Sigma_{i})$ has a random mean $\mu_{i}$ and a random covariance $\Sigma_{i}\in\{\Sigma_{1},\Sigma_{2}\}$. By randomly distributing the objects in each simulation, we ensure that the RL agent is reactive to the information gathered from the environment, rather than learning the specific locations of each object.
@@
A drone is flying over the field at a fixed altitude and has a camera with a FoV of size $F\times F\in\mathbb{N}^{2}$ facing downwards. The goal of the drone is to find all objects in the field as fast as possible. The drone can either start at the top-left or bottom-right part of the field, which is randomly selected. For experiment 1 - 3, the simulation terminates when all the objects are found, in experiment 4 we use a learned stop signal. Figure 2 shows two examples of the simulation environment.
@@
Figure 2: Two examples of the simulation environment with the field-of-view indicated by the blue rectangle around the drone, the flight path in blue, the detected objects in dark-grey and the not-yet detected objects in light-gray.
@@
The output of the detection network is simulated by creating a map of the objects that are visible within the FoV. Typical errors in the output of a detection network are False Positives (FPs), False Negatives (FNs) and positional errors. FP detections are simulated by adding $r_{\textrm{dt,fp}}\cdot F^{2}$ FP detections at a random location in the FoV where $r_{\textrm{dt,fp}}$ is the fraction FP detections with respect to the size of the FoV, $F$. FN detections are simulated by removing $r_{\textrm{dt,fn}}\cdot n_{\textrm{{fov}}}$ detections from the visible objects in the current FoV of the drone, were $r_{\textrm{dt,fn}}$ is the fraction of FN detections with respect to the number of objects visible in the current FoV, $n_{\textrm{fov}}$. Uncertainty in the position of the objects in the detection map is simulated by adding an offset drawn from the normal distribution $\mathcal{N}_{\textrm{dt,pos}}(0.0,\sigma)$, independently in both x and y directions. These errors are simulated independently for each detection.
@@
The prior knowledge is based on a map of ground truth locations of objects in the simulation. Inaccuracies in the prior knowledge are simulated by adding $r_{\textrm{pk,fp}}\cdot M^{2}$ false positives, where $r_{\textrm{pk,fp}}$ is the fraction of FPs in the prior knowledge with respect to the world size $M$. FNs in the prior knowledge are simulated by removing $r_{\textrm{pk,fn}}\cdot n$ objects from the ground truth map with the object locations where $r_{\textrm{pk,fn}}$ is the fraction of FNs with respect to the number of objects $n$. Same as for the detection network output, the location of the objects is altered by drawing an offset from the normal distribution $\mathcal{N}_{\textrm{pk,pos}}(0.0,\sigma)$. To simulate a reduction in resolution, the map is down-sampled using average pooling with a kernel size of $\left\lfloor\frac{M}{P}\right\rfloor$ where $P\times P$ is the resulting prior knowledge size. The prior knowledge is simulated once for each simulation.
@@ 2.2 Problem definition
To solve the described goal, the simulation is implemented as a Markov Decision Process (MDP). A MDP is described by a state-space $S$, an action-space $A$ and a reward $R$ (Sutton and Barto, 2018). In a state $s_{t}\in S$ at timestep $t$, an agent (the drone) performs an action $a_{t}\in A$ yielding a transition to state $s_{t+1}\in S$ and a reward $r_{t}\in R$.
@@
The state-space representation, $S$, is adapted from Theile et al. (2021), containing a global and local map representation and a movement budget scalar. The global map contains down-sampled information about the complete field (the prior knowledge) and the local map contains detailed information about the current FoV of the drone (representing the output of the detection network). Both global and local maps consist of three layers: a field-area layer, a layer with the locations of the already detected objects, and a third layer consisting of the prior knowledge for the global map and simulated detection network output for the local map. The field-area layer describes the area of the field (value of 0) and the area outside the field (value of 1). The layer with the already detected objects contains a value of 1 at the places where objects have been detected and 0 at all other locations. The movement budget scalar $b$ equals the remaining battery capacity and is calculated by $b=b_{\textrm{init}}-s\cdot b_{\textrm{step}}$, where $b_{\textrm{init}}$ is the initial battery level of the drone, $s$ the number of flight actions made in the field, and $b_{\textrm{step}}$ the battery usage of each step.
@@
Both the global and the local map are drone-centric, meaning that the drone is always in the middle of the map. Work by Theile et al. (2021) showed that centering the local and global map makes it possible to scale to larger fields. To do so, the global map is padded to a size of $(2M-1)\times(2M-1)$ by adding padding values of 0 for both the detected object and prior knowledge layer and values of 1 for the field-area layer (indicating that the drone is not allowed to fly outside the field). To decrease the size of the global map, the global map is down-sampled using average pooling with kernel size $g_{\textrm{global}}$ resulting in a global map size of $\frac{2M-1}{g_{\textrm{global}}}\times\frac{2M-1}{g_{\textrm{global}}}$. The local map is padded to a size of $F\times F$ by adding padding values of 0 for both the detected object map and output of the detection network, and values of 1 for the field-area layer when the FoV partly extends outside the field. An example of the global and local map is shown in Figure 3.
@@
Figure 3: Example of the local (a) and global (b) map. Red indicates the already detected objects, green the area outside the field and blue the simulated output of a detection network for the local map and the prior knowledge for the global map. Note that purple is a combination red and blue.
@@
The action-space, $A$, contains the following actions: fly north, fly south, fly east and fly west. Each fly action moves the drone one grid cell in the associated direction. The drone cannot move into the area outside the field. For experiment 4, we added another ’land’ action, that terminates the search.
@@
The default parameters for the simulation are given in Table 1. The simulation is implemented using the OpenAI Gym API and made available on GitHub [^1] https://github.com/wur-abe/rl_drone_object_search.
@@
$\mathcal{N}_{\textrm{obj}}(\mu,\sigma)$ $\mathcal{N}(100,30)$ Number of objects
@@
$\Sigma_{1}$ $[\begin{smallmatrix}5&8\\ 8&20\end{smallmatrix}]$ Covariance distribution 1
$\Sigma_{2}$ $[\begin{smallmatrix}20&0\\ 0&\end{smallmatrix}]$ Covariance distribution 2
@@
$r_{\textrm{step}}$0.2 Step reward

### v2

# UAV-based path planning for efficient localization of non-uniformly distributed weeds using prior knowledge: A reinforcement-learning approach

@@ Abstract
UAVs are becoming popular in agriculture, however, they usually use time-consuming row-by-row flight paths. This paper presents a deep-reinforcement-learning-based approach for path planning to efficiently localize weeds in agricultural fields using UAVs with minimal flight-path length. The method combines prior knowledge about the field containing uncertain, low-resolution weed locations with in-flight weed detections. The search policy was learned using deep Q-learning. We trained the agent in simulation, allowing a thorough evaluation of the weed distribution, typical errors in the perception system, prior knowledge, and different stopping criteria on the planner’s performance. When weeds were non-uniformly distributed over the field, the agent found them faster than a row-by-row path, showing its capability to learn and exploit the weed distribution. Detection errors and prior knowledge quality had a minor effect on the performance, indicating that the learned search policy was robust to detection errors and did not need detailed prior knowledge. The agent also learned to terminate the search. To test the transferability of the learned policy to a real-world scenario, the planner was tested on real-world image data without further training, which showed a 66% shorter path compared to a row-by-row path at the cost of a 10% lower percentage of found weeds. Strengths and weaknesses of the planner for practical application are comprehensively discussed, and directions for further development are provided. Overall, it is concluded that the learned search policy can improve the efficiency of finding non-uniformly distributed weeds using a UAV and shows potential for use in agricultural practice.
@@ 1 Introduction
In recent years, Unmanned Aerial Vehicles (UAVs) have become increasingly popular for various applications in agriculture (Rejeb et al., 2022). Examples in agriculture are finding weeds or diseased plants in large arable fields (Albani et al., 2019; Chin et al., 2023), detecting cattle in pastures (Rivas et al., 2018; Liu et al., 2021), and blossom detection in orchards (Zhang et al., 2023). Generally, these applications have a main task: to find objects of interest in an area larger than the field-of-view (FoV) of the UAV, which requires the UAV to fly over the field. A limitation of using UAVs for these applications is their limited battery capacity (Rejeb et al., 2022; Gugan and Haque, 2023). Therefore, it is important to find a path between all the objects of interest that minimizes the UAV’s flight time in order to increase the area that can be inspected in a single flight.
@@
When these objects are uniformly distributed over the area, a coverage path planner covering the entire search area is suitable and efficient. However, in applications where the objects are non-uniformly distributed, it may be more efficient to use a search policy that searches for objects instead of covering the whole area. A task such as the detection of weeds is an example of an application with non-uniformly distributed objects, since some weed species occur in distinct patches in a field (Dessaint et al., 1991; Cardina et al., 1997; Colbach et al., 2000). A lot of work is done on weed detection using drone images (Xu et al., 2023; Anul Haq, 2022; Pei et al., 2022); however, they all follow predefined, row-by-row, flight paths.
@@
Over the last years, Reinforcement Learning (RL) has gained more attention in path planning for both mobile robots (Yu et al., 2020; Gao et al., 2020; Niroui et al., 2019) as well as UAVs (Azar et al., 2021; Tu and Juang, 2023). These RL-based methods can learn search strategies through interaction with the environment by maximizing the information gathered in an environment (Lodel et al., 2022). RL offers several advantages for path planning: it does not need a detailed map of obstacles in the environment, can deal with noisy sensor information (Gao et al., 2020), and can learn spatial relations between objects of interest in the environment. Once trained, an RL learned policy makes sequential decisions about the direction of the robot or drone with relatively short calculation times, as compared to traditional methods like Dijkstra’s algorithm, A*, and D* (Tu and Juang, 2023). This allows for online path planning, eliminating the need to calculate the complete path in advance. Using RL, it is possible to plan paths with less or uncertain prior knowledge about the environment (Yu et al., 2020; Gugan and Haque, 2023), thereby making path planning more flexible and reactive to a changing environment (Popović et al., 2024).
Several works have studied path planning using RL. Panov et al. (2018) and Yu et al. (2020) showed that a trained RL network is able to generate a policy to move an agent in a grid-based world to a target location while avoiding obstacles. Tang et al. (2024) and Chronis et al. (2023) showed that an RL-learned path planner yielded shorter paths and required less computational time than traditional methods like A* for an environment with static obstacles. Castro et al. (2023) used RL to avoid obstacles while executing a flight path for inspecting fly traps in orchards. Westheider et al. (2023) showed an RL-based path planner for cooperative multi-UAV monitoring of terrains, and showed the applicability of the method by discovering temperature hotspots on a real-world thermal dataset of a crop field.
A novel neural network architecture for coverage path-planning and object search was introduced by Theile et al. (2020, 2021), combining detailed high-resolution local information with low-resolution global information. They show that this representation is efficient for UAV path planning to gather data from Internet-of-Things devices within a simulated city. Such representation can also be useful for a weed detection application using UAVs, where low-resolution global information about the field may be available on forehand, and local high-resolution information about the UAV’s field-of-view is available during flight. However, their approach assumes full prior knowledge about the location of the target devices and the obstacles, which is, in many agricultural applications, uncertain or even absent. For example, the location of weeds in a field may not be exactly known beforehand, however, some prior knowledge can be derived from the locations of the weeds during the previous year (Colbach et al., 2000; Dessaint et al., 1991). Uncertainty in prior knowledge also makes it more difficult to determine when to terminate the search task. When having full prior knowledge, there is enough information to stop the search task when all objects are found. However, when the prior knowledge becomes more sparse and unreliable, the number of objects might be unknown. This is especially true for weed detection, as the exact number of weeds in a field is unknown. This makes the end of the search task undefined. Yang et al. (2018) and Druon et al. (2020) used a learned stopping action, where the agent could decide to stop searching when it was not profitable anymore. Building on the idea of combining high-resolution local information and low-resolution global information in a single RL network architecture from Theile et al. (2021), we developed and evaluated a reinforcement-learning-based UAV path planner that can deal with this uncertain prior knowledge, to localize non-uniformly distributed weeds in a field and to decide when to terminate the search.
@@
We propose an RL-based path planner that works on a higher abstraction level, getting a local weed map extracted from a perception module based on a camera image, and providing actions in terms of flight directions that would be executed by a drone’s flight controller. Figure 1 illustrates how the resulting simulation-trained policy can be used in the real world to control the drone. The drone’s camera image, containing variations in the weeds’ appearance due to, e.g., lightning conditions and natural variation, is converted to a local weed map using an object detection network. The resulting weed map, combined with potential, uncertain weed locations derived from prior knowledge, is used as input to the RL policy. The policy determines the action with respect to the highest expected reward. The discrete output actions (fly north, east, west, south) are then translated to motor commands by the drone’s flight controller. The advantage of this abstraction is that it allows us to train the RL agent in simulation by simulating the detection network’s output and the prior knowledge, and to run the trained agent on real image data.
@@
Figure 1: High-level drone control using Reinforcement Learning (RL) in combination with a local weed map from a perception system and some potential weed locations from the prior knowledge. The output of the detection network is used together with prior knowledge as input for the RL policy, the resulting discrete actions are then translated by the drone’s flight controller to motor commands.
Simulation-based training is necessary because training an RL agent requires huge amounts of data, as it needs to explore an enormous state-action space (Gugan and Haque, 2023). For our use case, training an RL agent to localize weeds in a field would need the agent to collect a very large amount of training data containing a large variety of weed locations in the field to allow it to explore different policies. As this is infeasible to achieve in the real world, we trained the agent using simulations, as is commonly done in literature, e.g. Azar et al. (2021); Gao et al. (2020); Theile et al. (2020, 2021); Panov et al. (2018); Yu et al. (2020). An additional benefit of a simulation is that we can investigate the effect of different weed distributions, detection inaccuracies, and prior knowledge uncertainty in greater detail. It also allows for many more repetitions than would be feasible in real-world experiments and thereby shows the practical potential of such an RL-based path planner for localizing weeds in a field.
@@
However, when training in simulation, it is important that the abstract simulation includes typical errors in a perception model and a flight controller. The typical errors in translating the image into weed coordinates are false positive and false negative detections, and inaccuracies in the location of the detections. These are included in the simulation as explained in the materials section. Furthermore, the simulation makes some assumptions about the system: (a) it is assumed the drone is able to accurately execute discrete flight actions, such as fly one meter in a specific direction, (b) a maximum field size based on the number of grid-cells in the simulation and the area of each cell, (c) prior knowledge of the field shape and boundaries, (d) a trained detection network and (e) some potential, uncertain locations of weeds. The impact of all made assumptions is discussed at length in the discussion. To demonstrate the practical potential of learning RL policies in an abstract simulation, we show the performance of the simulation-learned RL policy on real-world image data.
@@
The objective of this paper is to develop and evaluate an RL-based method for path planning to efficiently localize weeds using UAVs with a minimal path length. The novelty of our method is the abstract representation of the environment, which allows for a real-world weed detection application in combination with an object detection network. Despite being mainly a simulation study, we provide an important contribution to a real-world application by analyzing the effects of the spatial distribution of the weeds, typical errors in detection network output, and typical errors associated with prior knowledge. To show the transferability of the simulation-learned search policy to a real-world application, we also evaluate its performance on real-world image data without further training. Specifically, we study (1) the effect of the distribution of the weeds, (2) the effect of detection errors, (3) the effect of the quality of prior knowledge, (4) the effect of different stopping criteria, and finally, (5) the performance of the RL-based path planner on real-world data.
@@ 2 Materials and Methods
In this chapter, we describe the simulation environment, problem statement, the used RL method, and the experiments.
@@ 2.1 Simulation environment
The field is simulated as a square grid of $M\times M\in\mathbb{N}^{2}$ grid cells, where $\mathbb{N}$ is the set of natural numbers. The real-world size of each grid cell depends on the required spatial resolution, which is not part of the simulation environment. In this field, the number of weeds, $n$ is drawn from a normal distribution $\mathcal{N}_{\textrm{obj}}(\mu,\sigma)$. These weeds are then distributed according to $k$ different multivariate Gaussian distributions, where $k$ is drawn from $\mathcal{N}_{\textrm{dist}}(\mu,\sigma)$. Each Gaussian distribution $\mathcal{N}(\mu_{i},\Sigma_{i})$ has a random mean $\mu_{i}$ and a random covariance $\Sigma_{i}\in\{\Sigma_{1},\Sigma_{2}\}$. By randomly distributing the weeds in each simulation, we ensure that the RL agent is reactive to the information gathered from the environment, rather than learning the specific locations of each weed. Additionally, by introducing randomness in the weed distributions, the RL agent can learn to respond to a large variety of different weed distributions through domain randomization.
@@
A drone is flying over the field at a fixed altitude and has a camera with a FoV of size $F\times F\in\mathbb{N}^{2}$ facing downward. The goal of the drone is to find all weeds in the field as fast as possible. The drone can start at the top-left or bottom-right part of the field, which is randomly selected. For experiments 1–3, the simulation terminates when all the weeds are found; in experiments 4 and 5, we use a learned stop signal. Figure 2 shows two examples of the simulation environment.
@@
Figure 2: Two examples of the simulation environment with the field-of-view indicated by the blue rectangle around the drone, the flight path in blue, the detected weeds in dark-gray and the not-yet detected weeds in light-gray.
@@
The output of the detection network is simulated by creating a map of the weeds that are visible within the FoV. Typical errors in the output of a detection network are False Positives (FPs), False Negatives (FNs), and positional errors. FP detections are simulated by adding $r_{\textrm{dt,fp}}\cdot F^{2}$ FP detections at a random location in the FoV, where $r_{\textrm{dt,fp}}$ is the fraction of FP detections with respect to the size of the FoV, $F$. FN detections are simulated by removing $r_{\textrm{dt,fn}}\cdot n_{\textrm{{fov}}}$ detections from the visible weeds in the current FoV of the drone, where $r_{\textrm{dt,fn}}$ is the fraction of FN detections with respect to the number of weeds visible in the current FoV, $n_{\textrm{fov}}$. Uncertainty in the position of the weeds in the detection map is simulated by adding an offset drawn from the normal distribution $\mathcal{N}_{\textrm{dt,pos}}(0.0,\sigma)$, independently in both x and y directions. These errors are simulated independently for each detection.
@@
To simulate prior knowledge, inaccuracies are added to the map of ground-truth locations of weeds. These inaccuracies are added using a similar approach as described above. False positives in the prior knowledge are added by adding $r_{\textrm{pk,fp}}\cdot M^{2}$ weeds to the ground-truth map, where $r_{\textrm{pk,fp}}$ is the fraction of FPs in the prior knowledge with respect to the field size $M$. FNs in the prior knowledge map are simulated by removing $r_{\textrm{pk,fn}}\cdot n$ weeds from the ground-truth map with the weed locations where $r_{\textrm{pk,fn}}$ is the fraction of FNs with respect to the number of weeds $n$. Same as for the detection network output, the location of the weeds is altered by drawing an offset from the normal distribution $\mathcal{N}_{\textrm{pk,pos}}(0.0,\sigma)$. To simulate a reduction in resolution, the map is down-sampled using average pooling with a kernel size of $\left\lfloor\frac{M}{P}\right\rfloor$ where $P\times P$ is the resulting prior knowledge size. The prior knowledge is simulated once for each simulation.
@@ 2.2 Problem definition
To solve the described goal, the simulation is implemented as a Markov Decision Process (MDP). An MDP is described by a state-space $S$, an action-space $A$, and a reward $R$ (Sutton and Barto, 2018). In a state $s_{t}\in S$ at timestep $t$, an agent (the drone) performs an action $a_{t}\in A$ yielding a transition to state $s_{t+1}\in S$ and a reward $r_{t}\in R$.
@@
The state-space representation, $S$, is adapted from Theile et al. (2021) and uses the same structure, containing a global and local map representation and a movement budget scalar. Maintaining the same structural representation, we encode different information in the global and local map: the global map contains down-sampled information about the complete field (the prior knowledge), and the local map contains detailed information about the current FoV of the drone (representing the output of the detection network). Both global and local maps consist of three layers: a field-area layer, a layer with the locations of the already detected weeds, and a third layer consisting of the prior knowledge for the global map and simulated detection network output for the local map. The field-area layer describes the area of the field (value of 0) and the area outside the field (value of 1). The layer with the already detected weeds contains a value of 1 at the places where weeds have been detected and 0 at all other locations. The movement budget scalar $b$ equals the remaining battery capacity and is calculated by $b=b_{\textrm{init}}-s\cdot b_{\textrm{step}}$, where $b_{\textrm{init}}$ is the initial battery level of the drone, $s$ the number of flight actions made in the field, and $b_{\textrm{step}}$ the battery usage of each step.
@@
Both the global and the local map are drone-centric, meaning that the drone is always in the middle of the map. Work by Theile et al. (2021) showed that centering the local and global map makes it possible to scale to larger fields. To do so, the global map is padded to a size of $(2M-1)\times(2M-1)$ by adding padding values of 0 for both the detected weed and prior knowledge layer and values of 1 for the field-area layer (indicating that the drone is not allowed to fly outside the field). To decrease the size of the global map, the global map is down-sampled using average pooling with kernel size $g_{\textrm{global}}$ resulting in a global map size of $\frac{2M-1}{g_{\textrm{global}}}\times\frac{2M-1}{g_{\textrm{global}}}$. The local map is padded to a size of $F\times F$ by adding padding values of 0 for both the detected weed map and output of the detection network, and values of 1 for the field-area layer when the FoV partly extends outside the field. An example of the global and local map is shown in Figure 3.
@@
Figure 3: Example of the local (a) and global (b) map. Red indicates the already detected weeds, green the area outside the field and blue the simulated output of a detection network for the local map and the prior knowledge for the global map. Note that purple is a combination of red and blue.
@@
The action-space, $A$, contains the following actions: fly north, fly south, fly east, and fly west. Each fly action moves the drone one grid cell in the associated direction. The drone cannot move into the area outside the field. For experiments 4 and 5, we added another ’land’ action that terminates the search.
@@
The default parameters for the simulation are given in Table 1. The simulation is implemented using the OpenAI Gym API and is published on GitHub [^1] https://github.com/wur-abe/rl_drone_object_search.
@@
$\mathcal{N}_{\textrm{obj}}(\mu,\sigma)$ $\mathcal{N}(100,30)$ Number of weeds
@@
$\Sigma_{1}$ $[\begin{smallmatrix}5&8\\ 8&15\end{smallmatrix}]$ Covariance distribution 1
$\Sigma_{2}$ $[\begin{smallmatrix}15&0\\ 0&5&\end{smallmatrix}]$ Covariance distribution 2
@@
$r_{\textrm{step}}$0.5 Step reward